OpenStudy (anonymous):

If a 7% saline solution and a 4% saline solution are mixed to make 500 milliliters of a 5% saline solution, how much of each solution, to the nearest milliliter, is needed?

OpenStudy (anonymous):

Let x be the number of milliliters of the 7% saline solution and y be the number of milliliters of the 4% saline solution. They need to add up to 500 milliliters total, so x + y = 500. The amount of salt in the x milliliter 7% solution is 0.07x milliliters. The amount of salt in the y milliliter 4% solution is 0.04y milliliters. Together, they need to add up to 5% of the total 500 milliliters, which is 0.05(500) = 25 milliliters, so 0.07x + 0.04y = 25. Solve the system by substitution. \[x + y = 500 \] \[x = 500 - y\] \[0.07x + 0.04y = 25\] (substitute 500 - y for x) \[0.07(500 - y) + 0.04y = 25\] \[35 - 0.07y + 0.04y = 25\] \[-0.03y + 35 = 25\] \[-0.03y = -10\] \[y = 333.333...\] y = about 333 \[x = 500 - y = 500 - 333 = 167\] ANSWER: 167 milliliters of the 7% solution and 333 milliliters of the 4% solution (to the nearest milliliter)